What is the angular speed in rad/s for a point on the surface of Mars due to its rotation on its axis? (radius=3400km, day =24.6 hrs)?

points will go to the clearest explination. thanks
Answer #1

Time Period of rotation is equal to = T=2\pi /(\omega) Time is T= 24.6 hours = 24.6*60*60 s = 88560 s

so angular velocity is w= (2*pi)/T = 7.091*10^(-5) rad/s